# A theoretical analysis on the resolution of parametric PDEs via Neural Networks designed with Strassen algorithm

### v2

@@ 2.2 Strassen Algorithm for Square Matrices
In this section, we recall the Strassen algorithm, which was proposed in [Str69]. First, we define the Strassen algorithm in matrices belonging to $\mathbb{R}^{2^{k}\times 2^{k}}$. Then, we define it in matrices in $\mathbb{R}^{n\times n}$ for arbitrary values $n\in\mathbb{N}$.
@@ 2.3 Matrix Neural Networks
We now define the notions of Matrix Neural Network, concatenation, and parallelization. This is based on the definition introduced in [PV18], but adapted for Neural Networks whose entries are matrices. We do that for the sake of simplifying the notation.
@@
Let $L,N^{0,1},N^{0,2}\in\mathbb{N}$. A Matrix Neural Network $\Phi$ with input dimensions $\operatorname{dim_{in}}\Phi\coloneq N^{0,1}\times N^{0,2}$ and $L$ layers is a sequence of triplets of tensors, matrices and activation functions:
$\Phi=((\mathcal{L}^{1},C^{1},\alpha^{1}),(\mathcal{L}^{2},C^{2},\alpha^{2}),\dots,(\mathcal{L}^{L},C^{L},\alpha^{L}))$
where $N^{1,1},N^{1,2},\dots,N^{L,1},N^{L,2}\in\mathbb{N}$, and where each $\mathcal{L}^{\ell}$ is a
$(N^{\ell,1}\times N^{\ell,2})\times(N^{\ell-1,1}\times N^{\ell-1,2})$
tensor, $C^{\ell}$ is a $N^{\ell,1}\times N^{\ell,2}$ matrix and
@@
for all $\ell=1,\dots,L$.
If $\Phi$ is a Matrix Neural Network as above and if $\varrho:\mathbb{R}\longrightarrow\mathbb{R}$, then we say that $\Phi$ is a $\varrho$-MNN if $\alpha^{\ell}_{i,j}\in\left\{\operatorname{id},\varrho\right\}$ for $i=1,\dots,N^{\ell,1}$, $j=1,\dots,N^{\ell,2}$, $\ell=1,\dots,L-1$, and $\alpha^{L}_{i,j}=\operatorname{id}$ for all $i=1,\dots,N^{L,1}$ and $j=1,\dots,N^{L,2}$.
Moreover, if $\Phi$ is a Matrix Neural Network as above, we define the realization of $\Phi$ as the map $R(\Phi):\mathbb{R}^{N^{0,1}\times N^{0,2}}\longrightarrow\mathbb{R}^{N^{L,1}\times N^{L,2}}$ such that
$R(\Phi)(A)=A^{L},$
where $A^{L}$ results from the following scheme:
$A^{0}$ :-$\coloneq A,:$
$A^{\ell}$ :-$\coloneq\alpha^{\ell}(\mathcal{L}^{\ell}A^{\ell-1}+C^{\ell}),\quad\text{ for }\ell=1,\dots,L.:$
We call $L=L(\Phi)$ the number of layers. For $\ell\leq K$ we call
$M_{\ell}(\Phi)\coloneq\operatorname{n_{W}}\left(\mathcal{L}^{\ell}\right)+\operatorname{n_{W}}\left(C^{\ell}\right)$
the number of weights in the $\ell$-th layer and we define
$M(\Phi)\coloneq\sum_{\ell=1}^{L}M^{\ell}(\Phi),$
which we call the number of weights of $\Phi$. Moreover, we refer to
$\operatorname{dim_{out}}\Phi\coloneq N^{L,1}\times N^{L,2}$
as the output dimension of $\Phi$.
Finally, given a previously defined Matrix Neural Network $\Phi$, in order to work with its layers, we write $((\mathcal{L}^{1},B^{1},\alpha^{1}),\dots,(\mathcal{L}^{L},B^{L},\alpha^{L}))\coloneq\Phi$ to specify which and in what order tensors, matrices, and activation functions conform to the Matrix Neural Network.
@@
Classical Neural Networks (if we consider them defined as in [PV18]) have the same structure as a Matrix Neural Networks: they are a sequence of triples of matrices, vectors, and activation functions. Their realization, number of layers, and number of weights are defined equivalently for these Neural Networks. In particular, any classical Neural Network can be viewed as a Matrix Neural Network by considering $N^{\ell,2}=1$ for all $\ell=1,\dots,N$. Conversely, using vectorization procedures, since the matrix space $\mathbb{R}^{n\times m}$ is isomorphic to the vector space $\mathbb{R}^{nm}$, all the Matrix Neural Networks can be implemented as classical Neural Networks (with the same length, number of weights, and final error). Therefore, we can view both notions as equivalent.
One important notion that we are going to use is the parallelization of Neural Networks:
Definition 2.9.
Let $L_{1},L_{2}\in\mathbb{N}$ and let
$\Phi^{1}$ $=\left((\mathcal{L}^{1,1},C^{1,1},\alpha^{1}),\dots,(\mathcal{L}^{1,L_{1}},C^{1,L_{1}},\alpha^{1,L_{1}})\right),$
$\Phi^{2}$ $=\left((\mathcal{L}^{2,1},C^{2,1},\alpha^{1}),\dots,(\mathcal{L}^{2,L_{2}},C^{2,L_{2}},\alpha^{2,L_{2}})\right)$
be two Matrix Neural Networks such that the input layer of $\Phi^{1}$ has the same dimension as the output layer of $\Phi^{2}$. Then, $\Phi^{1}\circ\Phi^{2}$ denotes the following $L_{1}+L_{2}$ layer Neural Network:
$\Phi^{1}\circ\Phi^{2}\coloneq\left((\mathcal{L}^{2,1},C^{2,1},\alpha^{2,1}),\dots,(\mathcal{L}^{2,L_{2}},C^{2,L_{2}},\alpha^{2,L_{2}}),\right.\\ \left.(\mathcal{L}^{1,1},C^{1,1},\alpha^{1,1}),\dots,(\mathcal{L}^{1,L_{1}},C^{1,L_{1}},\alpha^{1,L_{1}})\right)$
We call $\Phi^{1}\circ\Phi^{2}$ the concatenation of $\Phi^{1}$ and $\Phi^{2}$.
Remark 2.10.
In the literature, this concatenation is known as a sparse concatenation and is denoted by $\odot$. This difference comes from the fact that there is another kind of concatenation denoted by $\bullet$: given two $\varrho$-Matrix Neural Networks
$\Phi^{1}$ $=\left((\mathcal{L}^{1,1},C^{1,1},\alpha^{1}),\dots,(\mathcal{L}^{1,L_{1}},C^{1,L_{1}},\alpha^{1,L_{1}})\right),$
$\Phi^{2}$ $=\left((\mathcal{L}^{2,1},C^{2,1},\alpha^{1}),\dots,(\mathcal{L}^{2,L_{2}},C^{2,L_{2}},\alpha^{2,L_{2}})\right),$
such that $\operatorname{dim_{in}}\Phi^{1}=\operatorname{dim_{out}}\Phi^{2}$, the $\varrho$-Matrix Neural Network $\Phi^{1}\bullet\Phi^{2}$ is defined as:
$\Phi^{1}\bullet\Phi^{2}\coloneq\left((\mathcal{L}^{2,1},C^{2,1},\alpha^{2,1}),\dots,(\mathcal{L}^{2,L_{2}-1},C^{2,L_{2}-1},\alpha^{2,L_{2}-1}),\right.\\ \left.(\mathcal{L}^{1,1}\mathcal{L}^{2,L_{2}},\mathcal{L}^{1,1}C^{2,L_{2}}+C^{1,1},\alpha^{1,1}),\dots,(\mathcal{L}^{1,L_{1}},C^{1,L_{1}},\alpha^{1,L_{1}})\right).$
This $\varrho$-MNN satisfies again the property $R\left(\Phi^{1}\bullet\Phi^{2}\right)=R\left(\Phi^{1}\right)\circ R\left(\Phi^{2}\right)$. Even though this definition of concatenation produces Matrix Neural Networks with fewer layers, the estimation on the number of weights generally gets worse. This leads us to not use this definition.
Note that if $\Phi^{1}$ and $\Phi^{2}$ are a $\varrho$-MNN, then $\Phi^{1}\circ\Phi^{2}$ is also a $\varrho$-MNN too.
Definition 2.11.
Let $f^{i}:\mathbb{R}^{n_{i}\times n}\longrightarrow\mathbb{R}^{m_{i}\times m}$ for $i=1,2$ be two maps. Then, its tensor product $f^{1}\times f^{2}$ is the map defined as:
[…]
Remark 2.12.
It is easy to see that the tensor product of two tensors, viewed as maps between matrices spaces, is a tensor too.
Definition 2.13.
Let $L\in\mathbb{N}$ and let $\Phi^{1},\dots,\Phi^{k}$ be Neural Networks such that the input dimensions are shared and $\Phi^{i}=\left((\mathcal{L}^{i,1},C^{i,1},\alpha^{i,1}),\dots,(\mathcal{L}^{i,L},C^{i,L},\alpha^{i,L})\right)$. Then, we define the parallelization of $\Phi^{1},\dots,\Phi^{k}$ by:
$\mathcal{P}(\Phi^{1},\dots,\Phi^{k})\coloneq\left(\left(\mathcal{L}^{1,1}\times\dots\times\mathcal{L}^{k,1},\begin{pmatrix}C^{1,1}\\ \vdots\\ C^{k,1}\end{pmatrix},\alpha^{1,1}\times\dots\times\alpha^{k,1}\right),\dots,\right.\\ \left.\left(\mathcal{L}^{1,L}\times\dots\times\mathcal{L}^{k,L},\begin{pmatrix}C^{1,L}\\ \vdots\\ C^{k,L}\end{pmatrix},\alpha^{1,L}\times\dots\times\alpha^{k,L}\right)\right).$
@@
In order to implement the Strassen algorithm with $\varrho$-MM, with $\varrho$ fixed, we consider that we may multiply two numbers using Neural Networks. In particular, we suppose the following.
@@
An important example is ReLU, as we show in Section 6.
Another relevant example is $\varrho(t)=\max\left\{0,t\right\}^{2}$. In this case, the product is a $\varrho$-realization:
$xy=\frac{1}{4}\left(\varrho(x+y)+\varrho(-x-y)-\varrho(x-y)-\varrho(-x+y)\right)$
We have used that:
$\varrho(t)+\varrho(-t)=t^{2}$
and
$(x+y)^{2}-(x-y)^{2}=4xy.$
Thus, for $\varrho(t)=\max\left\{0,t\right\}^{2}$ there is a Neural Network $\Pi^{\varrho}_{0,\infty}$ satisfying:
$M(\Pi^{\varrho}_{0,\infty})=12,\quad L(\Pi^{\varrho}_{0,\infty})=2$
and computes the product exactly. This network is:
$\Pi^{\varrho}_{0,\infty}=\left(\left(\begin{pmatrix}1&1\\ -1&-1\\ 1&-1\\ -1&1\end{pmatrix},0,\varrho\times\varrho\times\varrho\times\varrho\right),\left(\frac{1}{4}\begin{pmatrix}1&1&-1&-1\end{pmatrix},0,\operatorname{id}\right)\right).$
3 Strassen Neural Networks
3.1 Strassen Algorithm for Square Matrices Whose Size is a Power of 2 Using Neural Networks
@@
$\Phi^{\operatorname{STR}_{2^{k},\varrho}}_{\varepsilon,K}=\Phi^{\operatorname{MIX}_{k}}\circ\Phi^{\operatorname{PAR}_{k,\varrho}}_{\varepsilon,K}\circ\Phi^{\operatorname{SPLIT}_{k}}.$
Let us show the following result with respect to the accuracy and size of the MNN that we have designed:
@@
Proof.
Let us suppose that Proposition 3.1 is true for $k-1$. Let us start by showing (3.4). Using the Strassen algorithm, we know that (2.2) is satisfied. Consequently, because of the triangular inequality, defining $P_{i}$ as in (2.3), it suffices to show that:
[…]
Let us show the estimate for $i=1$, as the other six estimates are analogous. Since $\left\lvert A\right\rvert_{\infty}\leq K$, then $\left\lvert A[1,1]+A[2,2]\right\rvert_{\infty}\leq 2K$. Similarly, since $\left\lvert B\right\rvert_{\infty}\leq K$, we have $\left\lvert B[1,1]+B[2,2]\right\rvert_{\infty}\leq 2K$. Thus, using the recursive hypothesis with $k-1$ we obtain that:
[…]
concluding the proof of (3.4).
Using the previous equality, (3.2) and (3.3), we obtain recursively:
$M\left(\Phi^{\operatorname{STR}_{2^{k},\varrho}}_{\varepsilon,K}\right)$ $=M\left(\Phi_{\varepsilon,K}^{\operatorname{PAR}_{k,\varrho}}\right)+M\left(\Phi^{\operatorname{MIX}_{k}}\right)+M\left(\Phi^{\operatorname{SPLIT}_{k}}\right)$ (3.8)
=$\cdots$= ⋯
We now calculate the sum using the fact that it is geometric:
$9\sum_{i=1}^{k}4^{i}\cdot 7^{k-i}=9\cdot 7^{k}\sum_{i=1}^{k}\left(\frac{4}{7}\right)^{i}=9\cdot 7^{k}\frac{4/7-(4/7)^{k+1}}{1-4/7}=12\left(7^{k}-4^{k}\right).$ (3.9)
Putting together (3.8) and (3.9), we obtain (3.5).
Thus, using induction, we conclude with (3.6). ∎
3.2 Strassen Algorithm for Arbitrary Square Matrices Using Neural Networks
Consider $m,n,p\in\mathbb{N}$. In order to multiply a matrix $A\in\mathbb{R}^{m\times n}$ by another matrix $B\in\mathbb{R}^{n\times p}$ using the Strassen algorithm implemented with Neural Networks, we are going to prolong the matrices to size $2^{k}\times 2^{k}$, for:
@@ 3.2 Strassen Algorithm for Arbitrary Square Matrices Using Neural Networks
$\Phi_{\varepsilon,K}^{\operatorname{STR}_{m,p,\varrho}^{n}}=\Phi^{\operatorname{SHR}_{n}}_{m,p}\circ\Phi_{\varepsilon,K}^{\operatorname{STR}_{2^{k_{m,p}^{n}},\varrho}}\circ\Phi^{\operatorname{EXT}_{n}}_{m,p}.$
Proposition 3.2.
@@
Then, for every $\varepsilon>0$,
@@
$M\left(\Phi_{\varepsilon,K}^{\operatorname{STR}_{m,p,\varrho}^{n}}\right)\leq 7\max\left\{m,n,p\right\}^{\log_{2}7}\left(M\left(\Pi^{\varrho}_{\frac{\varepsilon}{4\max\left\{m,n,p\right\}^{2}},2\max\left\{m,n,p\right\}K}\right)+12\right)$ (3.13)
$\phantom{aaaaaaaaaaaaa}-9\max\left\{m,n,p\right\}^{2},$
$L\left(\Phi_{\varepsilon,K}^{\operatorname{STR}_{m,p,\varrho}^{n}}\right)\leq L\left(\Pi_{\frac{\varepsilon}{4\max\left\{m,n,p\right\}^{2}},2\max\left\{m,n,p\right\}K}^{\varrho}\right)+2(\log_{2}(\max\left\{m,n,p\right\})+2).$ (3.14)
Proof.
Considering (3.4), we clearly have (3.12).
In addition, the number of weights can be obtained using (3.5), (3.11) and (3.10):
$M\left(\Phi_{\varepsilon,K}^{\operatorname{STR}_{m,p,\varrho}^{n}}\right)$ $=mp+M\left(\Phi^{\operatorname{STR}_{2^{k^{n}_{m,p}},\varrho}}_{\varepsilon,K}\right)+n(m+p)$
$=7^{k^{n}_{m,p}}\left(M\left(\Pi^{\varrho}_{\varepsilon/4^{k^{n}_{m,p}},2^{k^{n}_{m,p}}K}\right)+12\right)-12\cdot 4^{k^{n}_{m,p}}+mp+mn+np$
$\leq 7\max\left\{m,n,p\right\}^{\log_{2}7}\left(M\left(\Pi^{\varrho}_{\frac{\varepsilon}{4\max\left\{m,n,p\right\}^{2}},2\max\left\{m,n,p\right\}K}\right)+12\right)$
$-12\max\left\{m,n,p\right\}^{2}+mp+mn+np$
$\leq 7\max\left\{m,n,p\right\}^{\log_{2}7}\left(M\left(\Pi^{\varrho}_{\frac{\varepsilon}{4\max\left\{m,n,p\right\}^{2}},2\max\left\{m,n,p\right\}K}\right)+12\right)$
$-9\max\left\{m,n,p\right\}^{2}.$
To finish, regarding the length of the Neural Network, considering (3.6), (3.10), and that we have added two shallow Neural Networks:
$L\left(\Phi_{\varepsilon,K}^{\operatorname{STR}_{m,p,\varrho}^{n}}\right)$ $=L\left(\Phi_{m,p}^{\operatorname{SHR}_{n}}\right)+L\left(\Phi_{\varepsilon,K}^{\operatorname{STR}_{2^{k_{m,p}^{n}},\varrho}}\right)+L\left(\Phi_{m,p}^{\operatorname{EXT}_{n}}\right)$
$=L\left(\Pi_{\varepsilon/4^{k_{m,p}^{n}},2^{k_{m,p}^{n}}K}^{\varrho}\right)+2(k_{m,p}^{n}+1)$
$\leq L\left(\Pi_{\frac{\varepsilon}{4\max\left\{m,n,p\right\}^{2}},2\max\left\{m,n,p\right\}K}^{\varrho}\right)+2(\log_{2}(\max\left\{m,n,p\right\})+2).$
When we consider square matrices, we can extend the definition of the Neural Network $\Phi^{\operatorname{STR}_{n,\varrho}}_{\varepsilon,K}$ to $n\neq 2^{k}$ as follows:
$\Phi^{\operatorname{STR}_{n,\varrho}}_{\varepsilon,K}=\Phi_{n,n}^{\operatorname{SHR}_{n}}\circ\Phi_{\varepsilon,K}^{\operatorname{STR}_{2^{k},\varrho}}\circ\Phi_{n}^{\operatorname{EXT}^{*}},$ (3.15)
@@
for all $A,B\in\mathbb{R}^{n\times n}$. Note that we have the following equality:
$M\left(\Phi_{n}^{\operatorname{EXT}^{*}}\right)=2n^{2}.$
@@
When $n=2^{k}$, $R\left(\Phi^{SHR_{n}}_{n,n}\right)$ and $R\left(\Phi_{n}^{EXT^{*}}\right)$ are identity tensors, which is why we consider (3.15) a generalization. In that case, there is only a small difference in its length and number of weights, which is irrelevant for studying their asymptotic values.
Considering Proposition 3.2, we have that:
Theorem 3.4.
@@
Indeed, the resulting Matrix Neural Network is as in Proposition 3.2 but without transposing.
4 A Neural Network for Matrix Invertion

### v3

@@ 2.2 Strassen Algorithm for Square Matrices
In this section, we recall the Strassen algorithm, which was proposed in [strassen1969gaussian]. First, we define the Strassen algorithm in matrices belonging to $\mathbb{R}^{2^{k}\times 2^{k}}$. Then, we define it in matrices in $\mathbb{R}^{n\times n}$ for arbitrary values $n\in\mathbb{N}$.
@@ 2.3 Matrix Neural Networks
We now define the notions of Matrix Neural Network, concatenation, and parallelization. This is based on the definition introduced in [petersen2018optimal] and [gribonval2022approximation], but adapted for Neural Networks whose entries are matrices. We do this for the sake of simplifying the notation.
@@
Let $\varrho:\mathbb{R}\longrightarrow\mathbb{R}$. A $\varrho$-Matrix Neural Network ($\varrho$-MNN from now on) is a tuple $((\mathcal{L}^{1},C^{1},\alpha^{1}),\dots,(\mathcal{L}^{L},C^{L},\alpha^{L}))$ where each $\mathcal{L}^{\ell}$ is a $(N^{\ell,1}\times N^{\ell,2})\times(N^{\ell-1,1}\times N^{\ell-1,2})$ tensor, $C^{\ell}$ is a $N^{\ell,1}\times N^{\ell,2}$ matrix and
@@
satisfying $\alpha^{\ell}_{i,j}\in\left\{\operatorname{id},\varrho\right\}$ if $\ell<L$ and $\alpha^{L}=\operatorname{id}$.
@@
Multilayer Perceptron (if we consider them defined as in [petersen2018optimal]) have the same structure as a Matrix Neural Networks: they are a sequence of triples of matrices, vectors, and activation functions. Their realization, number of layers, and number of weights are defined equivalently for these Neural Networks. In particular, any Multilayer Perceptron can be viewed as a Matrix Neural Network by considering $N^{\ell,2}=1$ for all $\ell=1,\dots,N$. Conversely, using vectorization procedures, since the matrix space $\mathbb{R}^{n\times m}$ is isomorphic to the vector space $\mathbb{R}^{nm}$, all the Matrix Neural Networks can be implemented as classical Neural Networks (with the same length, number of weights, and final error). Therefore, we can view both notions as equivalent.
We can directly extend for a $\varrho$-MNN $\Phi$ the definitions given in [kutyniok2022theoretical] of its input and output dimensions $\operatorname{dim_{in}}(\Phi),\operatorname{dim_{out}}(\Phi)$, number of layers $L(\Phi)$, number of weights at the layer $\ell$ $M_{\ell}(\Phi)$, number of weights $M(\Phi)$ and realization $R(\Phi)$. We omit their formal definitions for the sake of brevity as they are standard notions.
For more than one $\varrho$-MNN $\Phi^{1},\dots,\Phi^{k}$, we can again extend the ideas of their sparse concatenation $\Phi^{1}\odot\Phi^{2}$ and their parallelization $\mathcal{P}(\Phi^{1},\dots,\Phi^{k})$ as defined in, for example, [petersen2018optimal] and [kutyniok2022theoretical]. In these papers, they focused in the activation function $\operatorname{ReLU}$ and these definitions are written with that in mind, but, allowing components of the activation functions to be the identity solves this technicalities. The whole point of introducing these concepts is that they satisfy the following:
$R(\Phi^{1}\odot\Phi^{2})=R(\Phi^{1})\circ R(\Phi^{2})$
and
$R(\mathcal{P}(\Phi^{1},\dots,\Phi^{k}))\begin{pmatrix}A_{1}\\ \vdots\\ A_{k}\end{pmatrix}=\begin{pmatrix}R(\Phi_{1})(A_{1})\\ \vdots\\ R(\Phi_{k})(A_{k})\end{pmatrix}$,
while keeping track in its number of weights and layers.
@@
In order to implement the Strassen algorithm with $\varrho$-MNN, with $\varrho$ fixed, we consider that we may multiply two numbers using Neural Networks. In particular, we suppose the following.
@@
This is easily satisfied by the multiple versions of the Universal Approximation Theorem, for example, that in [cybenko1989approximation]. Actually, Hypothesis 1 is equivalent to these results: the product approximation unlocks the polynomial approximation and therefore, the approximation of continuous functions.
Remark 2.5.
For the particular case $\varrho=\operatorname{ReLU}=t\mapsto\max\left\{t,0\right\}$, $\Pi_{\varepsilon,K}^{\operatorname{ReLU}}$ is constructed in [elbrachter2022dnn], which, by a minor modification using the equality
$4xy=(x+y)^{2}-(x-y)^{2}$
instead of
$2xy=(x+y)^{2}-(x^{2}+y^{2}),$
we can slightly improve the upper bounds on the number of weights:
$M\left(\Pi_{\varepsilon,K}^{\operatorname{ReLU}}\right)\leq\begin{cases}30\log_{2}K+15\log_{2}\frac{1}{\varepsilon}+25&\text{ if }\varepsilon<\frac{K^{2}}{2},\\ 12&\text{ if }\varepsilon\in\left[\frac{K^{2}}{2},K^{2}\right),\\ 0&\text{ if }\varepsilon\geq K^{2},\end{cases}$ (2.4)
and on the number of layers:
$L\left(\Pi_{\varepsilon,K}^{\operatorname{ReLU}}\right)\leq\begin{cases}\log_{2}K+\frac{1}{2}\log_{2}\frac{1}{\varepsilon}+\frac{5}{2}&\text{ if }\varepsilon<\frac{K^{2}}{2},\\ 2&\text{ if }\varepsilon\in\left[\frac{K^{2}}{2},K^{2}\right),\\ 1&\text{ if }\varepsilon\geq K^{2}.\end{cases}$ (2.5)
Remark 2.6.
An easy example is $\varrho=\operatorname{ReLU}^{2}$ where we can compute the product exactly. This is proven by just noticing that $\varrho(t)+\varrho(-t)=t^{2}$.
3 Main Results
3.1 Matrix Multiplication Approximation
@@
$\Phi^{\operatorname{STR}_{2^{k},\varrho}}_{\varepsilon,K}=\Phi^{\operatorname{MIX}_{k}}\odot\Phi^{\operatorname{PAR}_{k,\varrho}}_{\varepsilon,K}\odot\Phi^{\operatorname{SPLIT}_{k}}.$
Its properties regarding accuracy and number of parameters are summarized in the following proposition.
@@
For a detailed proof, see Appendix A.
Let us now extend this results to matrices of any shape. Consider $m,n,p\in\mathbb{N}$. In order to multiply a matrix $A\in\mathbb{R}^{m\times n}$ by another matrix $B\in\mathbb{R}^{n\times p}$ using the Strassen algorithm implemented with Neural Networks, we are going to prolong the matrices to size $2^{k}\times 2^{k}$, for:
@@
$\Phi_{\varepsilon,K}^{\operatorname{STR}_{m,p,\varrho}^{n}}=\Phi^{\operatorname{SHR}_{n}}_{m,p}\odot\Phi_{\varepsilon,K}^{\operatorname{STR}_{2^{k_{m,p}^{n}},\varrho}}\odot\Phi^{\operatorname{EXT}_{n}}_{m,p}.$
Corollary 3.2.
@@
Then, if we call $\gamma=\max\left\{m,n,p\right\}$, we have for every $\varepsilon>0$ the following:
@@
$M\left(\Phi_{\varepsilon,K}^{\operatorname{STR}_{m,p,\varrho}^{n}}\right)\leq 7\gamma^{\log_{2}7}\left(M\left(\Pi^{\varrho}_{\frac{\varepsilon}{4\gamma^{2}},2\gamma K}\right)+12\right)-9\gamma^{2},$ (3.10)
$L\left(\Phi_{\varepsilon,K}^{\operatorname{STR}_{m,p,\varrho}^{n}}\right)\leq L\left(\Pi_{\frac{\varepsilon}{4\gamma^{2}},2\gamma K}^{\varrho}\right)+2(\log_{2}\gamma+2).$ (3.11)
This result is a direct consequence of 3.1 and the definition of $\Phi^{\operatorname{STR}}$ and $\Phi^{\operatorname{EXT}}$.
Remark 3.3.
When considering $\varrho=\operatorname{ReLU}$, we get, using (2.4) and (2.5), the following asymptotic bounds:
$M\left(\Phi^{\operatorname{STR}_{m,p,\varrho}^{n}}_{\varepsilon,K}\right)$ $\leq C\gamma^{\log_{2}7}\left[\log_{2}\frac{1}{\varepsilon}+\log_{2}\gamma+\log_{2}\max\left\{1,K\right\}+1\right],$ (3.12)
$L\left(\Phi^{\operatorname{STR}_{m,p,\varrho}^{n}}_{\varepsilon,K}\right)$ $\leq C\left[\log_{2}\frac{1}{\varepsilon}+\log_{2}\gamma+\log_{2}\max\left\{1,K\right\}+1\right].$ (3.13)
Compared to the result presented in [kutyniok2022theoretical, Proposition 3.7], we reduce the dimension dependency of order 3 to $\log_{2}7$, as expected by applying Strassen’s algorithm.
For square matrices, to avoid using the transpose in the input, we only require to do a minor modification on the definition of the previous $\varrho$-MNN:
$\Phi^{\operatorname{STR}_{n,\varrho}}_{\varepsilon,K}=\Phi_{n,n}^{\operatorname{SHR}_{n}}\odot\Phi_{\varepsilon,K}^{\operatorname{STR}_{2^{k},\varrho}}\odot\Phi_{n}^{\operatorname{EXT}^{*}},$ (3.14)
@@
for all $A,B\in\mathbb{R}^{n\times n}$.
@@
When $n=2^{k}$, $R\left(\Phi^{\operatorname{SHR}_{n}}_{n,n}\right)$ and $R\left(\Phi_{n}^{\operatorname{EXT}^{*}}\right)$ are identity tensors, which is why we consider (3.14) an extension. In that case, there is only a small difference in its length and number of weights, which is irrelevant for studying their asymptotic values.
Considering Corollary 3.2, we have that:
Corollary 3.5.
@@
Indeed, the resulting Matrix Neural Network is as in Corollary 3.2 but without transposing.
3.2 Matrix Inversion Approximation